OpenStudy (anonymous):

(square root of three + 1)^2 Explain with work?

OpenStudy (anonymous):

(sqrt{3} + 1)^2 (sqrt{3} + 1)(sqrt{3} + 1) 3 + 2sqrt{3} + 1 4 + 2sqrt{3} 2(2 + sqrt{3})

OpenStudy (anonymous):

= 3 + 2sqrt3 + 1 = 4 + 2sqrt3

OpenStudy (anonymous):

\[(\sqrt{3} + 1)^2\]=3 + 1+ 2 * sqrt(3) = 4+ 2* 1.73 =7.46

OpenStudy (anonymous):

where i used the formula (a+b)^2=a^2+b^2+2*a*b

OpenStudy (anonymous):

Thanks Jimmyrep, ramkrishna & pk51.(: